OpenStudy (anonymous):

Calculate the average rate of change for the given function, from x = −3 to x = 7. x | F(x) ----- -3 | 50 3 | 10 7 | 0 Please help me with this, and please DO NOT direct answer this question. I want to know how to solve this question. Thank you :) *Would give you a medal award if you help me :D

OpenStudy (anonymous):

(0-50)/(7- (-3)) = -50/10 = -5

OpenStudy (anonymous):

Oh! Now, I got it! Thank you so much! You definitely make it very simple than the explanation from algebra II lol